OpenStudy (cody88):

are 10 8 6 measures of a right triangle?

OpenStudy (anonymous):

a²+b²=c²

OpenStudy (cody88):

wat yu mean by that

OpenStudy (anonymous):

that's the pathogrean theorem so... 64+36=100 which means that is a right triangle

OpenStudy (cody88):

thx how yu get those numbers tho

OpenStudy (anonymous):

So if you take the two smallest side and multiply them by themselves and then add them together Then multiply the largest number by itself. If they make a true statement than the triangle is a right triangle...Understand?

OpenStudy (anonymous):

a²+b²=c²

OpenStudy (cody88):

ok wat about 13 12 5

OpenStudy (anonymous):

ill walk you though it....5*5 =? 12*12=? 13*13=?

OpenStudy (cody88):

ok i got it thx

OpenStudy (anonymous):

no prob

OpenStudy (cody88):

so 17 14 and 6 are not right?

OpenStudy (anonymous):

36*196=

OpenStudy (cody88):

but i did it th other way

OpenStudy (anonymous):

36+196 do not equal 17² so no it isnt a right triangle

OpenStudy (anonymous):

6=a² 14=b² 17=c²

OpenStudy (anonymous):

all this is right now is substitution.

OpenStudy (cody88):

yea i get it thnx

OpenStudy (anonymous):

no problem

OpenStudy (cody88):

ok now i gotta find th leg i subtract right?

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (cody88):